# Language Guided Skill Discovery

### v1

@@ E.1 Observations
EEF quat Rotation of Robot arm’s end-effector 4
@@ E.2 Hyperparameters
Optimizer Adam(Kingma and Ba [2014])
Minibatch size 16384
@@
Appendix F Prompts used for each of the experiments
In this section, we share the prompts used for each of the experiments, and the examplary outputs from the LLM. One key detail is that we prompted LLM to finalize its sentence with a keyword of "Description:", so that we can parse the final description out of whole generated sentences.
F.1 Prompts for enforcing Ant agent to move North
@@
F.2 Prompts for enforcing Franka agent to push objects toward North
@@
F.3 Prompts for Franka agent manipulates object in general

### v2

@@ F.2.1 Observations
Gravity projection Vector indicates direction of the gravity 3
@@
EEF quat Rotation of the Robot arm’s end-effector 4
@@ F.2.2 Hyperparameters
Optimizer Adam(Kingma & Ba, 2014)
Minibatch size 32768(Ant) , 16384(Franka)
@@
Appendix G Prompts used for each of the experiments
In this section, we share the prompts used for each of the experiments, and the examplary outputs from the LLM. One key detail is that we prompted LLM to finalize its sentence with a keyword of ”Description:”, so that we can parse the final description out of whole generated sentences.
G.1 Prompts for enforcing Ant agent to move North
@@
G.2 Prompts for enforcing Franka agent to push objects toward North
@@
G.3 Prompts for Franka agent manipulates object in general